OpenStudy (anonymous):

Trig Identity please help!... (sin x - cos x)^2 = 1-sin 2x

OpenStudy (anonymous):

are you trying to prove this to be true???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Here is my best shot - I'm out of practice: (sin x - cos x)(sin x -cos x) = \[\sin ^{2}x - 2 (\sin x \cos x) +\cos^{2}x\]...(Pythagorean Identity)-->\[\sin ^{2}x + \cos ^{2} x =1\] (Double Angle formula) --> \[\sin 2 x = 2 \sin x \cos x\] - from there you can prove your equation to be true

OpenStudy (anonymous):

im looking.. it looks right

OpenStudy (anonymous):

THANK YOU, I really appreciate it!